OpenStudy (anonymous):

What is the value of w? 2(2w-6)=6(w-6)+14 W=?

OpenStudy (anonymous):

Simplifying 2(2w + -6) = 6(w + -6) + 14

OpenStudy (anonymous):

Reorder the terms: 2(-6 + 2w) = 6(w + -6) + 14 (-6 * 2 + 2w * 2) = 6(w + -6) + 14 (-12 + 4w) = 6(w + -6) + 14

OpenStudy (anonymous):

Reorder the terms: -12 + 4w = 6(-6 + w) + 14 -12 + 4w = (-6 * 6 + w * 6) + 14 -12 + 4w = (-36 + 6w) + 14

OpenStudy (anonymous):

Reorder the terms: -12 + 4w = -36 + 14 + 6w

OpenStudy (anonymous):

Combine like terms: -36 + 14 = -22 -12 + 4w = -22 + 6w

OpenStudy (anonymous):

Ya im not very good at math :P, sorry

OpenStudy (anonymous):

lol its ok

OpenStudy (anonymous):

Solving -12 + 4w = -22 + 6w

OpenStudy (anonymous):

there is still allot more.... continue?

OpenStudy (anonymous):

sure that would be nice XD

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

maybe ill catch on.

OpenStudy (anonymous):

Solving for variable 'w'

OpenStudy (anonymous):

Move all terms containing w to the left, all other terms to the right.

OpenStudy (anonymous):

Add '-6w' to each side of the equation. -12 + 4w + -6w = -22 + 6w + -6w

OpenStudy (anonymous):

so 6w and 2w?? or 2w and w-6?

OpenStudy (anonymous):

6w and 2w

OpenStudy (anonymous):

Combine like terms: 6w + -6w = 0 -12 + -2w = -22 + 0 -12 + -2w = -22

OpenStudy (anonymous):

still a little confused...

OpenStudy (anonymous):

im not done yet hold on...

OpenStudy (anonymous):

Add '12' to each side of the equation. -12 + 12 + -2w = -22 + 12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Combine like terms: -12 + 12 = 0 0 + -2w = -22 + 12 -2w = -22 + 12

OpenStudy (anonymous):

Combine like terms: -22 + 12 = -10 -2w = -10

OpenStudy (anonymous):

Divide each side by '-2'. w = 5

OpenStudy (anonymous):

Simplifying w = 5

OpenStudy (anonymous):

Thanks so much :D

OpenStudy (anonymous):

np :)